OpenStudy (sparklestaraa):

A system of equations is shown below: 6x = -7y + 11 5x + 2y = 13 What is the solution to this system of equations? (-3, 1) (-1, 3) (1, -3) (3, -1)

OpenStudy (sparklestaraa):

@jim_thompson5910

OpenStudy (jdoe0001):

if you solve the 1st equation for "y", what does it give you?

OpenStudy (jdoe0001):

or for "x", that'd work as well

OpenStudy (sparklestaraa):

-11?

OpenStudy (sparklestaraa):

@Luigi0210

OpenStudy (sparklestaraa):

@perl

OpenStudy (sparklestaraa):

@DanJS PLEASEEEEEE help pleaseeeeeeeee

OpenStudy (danjs):

6x + 7y = 11 (x 2) 5x + 2y = 13 (x -7) ------------- 12x + 14y = 22 -35x - 14 y = -91 ------------------ add them together,

OpenStudy (danjs):

-23x = -69

OpenStudy (danjs):

x=3 put that into one of the equations and get Y... done

OpenStudy (sparklestaraa):

Ohhhh so option A :D

OpenStudy (danjs):

each option is a point (x,y) and x=3